The pauling electronegativities of fluorine & chlorine are 4.0 & 3.0 respectively estimate the dipolemoment of chlorinemono fluorine(ClF).

Grade:12th pass

1 Answers

SURYA PRAKASH ROHIN
5 years ago
We can use an equation to estimate the dipole moment of a diatomic molecule from the Pauling electronegativities. 

u/D = deltaX = XF - XCl = 4.0 - 3.0 = 1.0 D 

Fluorine is more electronegative than chlorine, so the dipole is directed from fluorine to chlorine.